# Classification of curl forces for all space dimensions

### v2

@@ 3 Force decomposition
The decomposition $\omega=dH\omega+Hd\omega$ is therefore satisfied, with anriexact part being zero.
@@
$dH\omega=d(\frac{1}{2}(xy-yx))=0,$ (28)
@@
We can now proceed with applying the (non-unique) Frobenius decomposition. We need to select $\Gamma$ and $\Sigma$ such that the equation $d\Omega=\Gamma\wedge\Omega+\Sigma$ is satisfied.A straightforward approach is to first select an factor $e^{-g}$, which fixes $\Gamma$, and then determine the corresponding $\Sigma$.
If we select $e^{-g}=x$, then $g=-\ln(x)$ and $\Gamma=dg=\frac{-1}{x}dx$. This gives $\Sigma=d\Omega-\Gamma\wedge\Omega=3dx\wedge dy$. We can then compute the potentials: $h=H(e^{-g}\Omega)=H(x^{2}dy-xydx)=0$. Since $d\Gamma=0$, we have $\Sigma^{\prime}=\Sigma$, and $\eta=H(e^{-g}\Sigma)=H(3xdx\wedge dy)=x^{2}dy-xydx=x\Omega$.
For an another choice, let $e^{-g}=x^{2}$. Analogous calculations yield $g=-\ln(x^{2})$, $\Gamma=\frac{-2}{x}dx$, $\Gamma\wedge\Omega=-2dx\wedge dy$, and $\Sigma=\Sigma^{\prime}=4dx\wedge dy$. The potentials in this case are $h=0$ and $\eta=x^{3}dy-x^{2}ydx$.
Comparing these two selections of the $e^{-g}$ term demonstrates that the Frobenius decomposition is not fixed; the corresponding terms $\Gamma$ and $\Sigma$ can be adjusted to find a case that is easier to calculate.
@@
$\Omega=Hd\omega=H(2ydy\wedge dx)=\frac{2}{3}(y^{2}dx-xydy)=\omega-dH\omega.$ (32)
Note that both terms of the original form $\omega$ contribute to exact part. contribute to the final exact part. This shows that even if a component of a form is not exact, it can still contribute to the exact part of the total form. The potential for exact term is
@@
We therefore have $e^{g}=\frac{2}{3}xy^{2}$ and $h=\ln\left(\frac{x}{y}\right)$ and it is visible that it is a gradient recursive case.
@@ 4 Discussion
The exact part of the work form is associated with the conservative force, a classical result in mechanics. The antiexact part, as was noted in [19], can be thought of as the influence of an external agent (or a non-autonomous influence) from outside the system.
Its further decomposition using the Frobenius theorem extracts the integrable part, $e^{g}dh$, which can be interpreted in terms of inaccessibility, as in the Carathèodory interpretation of thermodynamics; see [26]. In the gradient recursive case, this is the only term comprising the antiexact part $\Omega$. It can be interpreted as a standard potential-driven force derived from $h$, but where the system’s coupling to this potential is modulated by a position-dependent scaling factor $e^{g}$. Physically, it could model phenomena like a charged particle moving in a magnetic field where the field strength varies in a specific way, or certain types of controlled robotic manipulations.
For the full and recursive case (but not for the gradient recursive case), the term $e^{g}\eta$ is also presents. It represents a more significant obstruction to integrability, as it cannot be written as the exterior derivative of a function multiplied by another function. This term represents the ”path-dependent core” of the force, which cannot be eliminated by a simple scaling factor. It can be present in driven systems where the external agent’s input is not derivable from any potential, scaled or otherwise. Note that this term is present in the full case (when torsion $\Sigma\neq 0$) or in the recursive case (when torsion vanishes, however when $\Gamma$ has antiexact part). It therefore measures the torsion and the antiexactness of $\Gamma$.
@@
The operator $d^{\nabla}$ bears a resemblance to the exterior covariant derivative on the trivial line bundle (or fiber set [19, 10]) $U\times\mathbb{R}\rightarrow U$. In general problem of parallel transport on a bundle, one provides the connection form $\Gamma$ and an external force $\Sigma$ to compute $\Omega$, as was presented in [19, 10]. However, for (6), we obtained the inverse problem: determine $\Gamma$ and $\Sigma$ for a fixed $\Omega$. This is a non-unique problem. Equation (35) also suggests that the antiexact force is a solution of a non-uniform parallel transport problem on a bundle with connection form $\Gamma$ and non-uniformity term $\Sigma$. As we infer in [19], the antiexact part $\Omega$ represents an external influence on the system, and equation (35) can be interpreted as an evolution of this external agent, seen through the force $\Omega$ with which it influence the system. Moreover, the curvature for this external system is given by [^5] In general the curvature of a connection form $\Gamma$ is given by $\Theta=d\Gamma+\Gamma\wedge\Gamma$. However, for a $\mathbb{R}$-valued form the term $\Gamma\wedge\Gamma$ vanishes. $\Theta=d\Gamma$.
@@ 5 Conclusions
A local decomposition for classical forces in a space of arbitrary dimension is proposed. This method employs the geometric decomposition on a star-shaped subset of a manifold, which distinguishes the exact part, associated with the gradient of a potential, from an antiexact part that generalizes the curl component of the force. The antiexact part is further resolved using the (non-unique) Frobenius decomposition into an integrable component and a remainder that cannot be represented as the exterior derivative of a function.
@@ Acknowledgments
The authors declare no conflict of interests.
@@
The work was partially funded by the EU Horizon Europe MSCA grant No 101183077. The author would like to thanks Michael V. Berry and Pragya Shukla for enlightening discussion on the nature of curl forces.

### v3

@@ 3.2 Examples
The decomposition $\omega=dH\omega+Hd\omega$ is therefore satisfied, with the antiexact part being zero.
@@
$dH\omega=d\left(\frac{1}{2}(xy-yx)\right)=0,$ (28)
@@
We can now proceed with applying the (non-unique) Frobenius decomposition. We need to select $\Gamma$ and $\Sigma$ such that the equation $d\Omega=\Gamma\wedge\Omega+\Sigma$ is satisfied. A straightforward approach is first to select a factor $e^{-\gamma}$ that fixes $\Gamma$ and then determine the corresponding $\Sigma$.
If we select $e^{-\gamma}=x$, then $\gamma=-\ln(x)$ and $\Gamma=d\gamma=\frac{-1}{x}dx$. This gives $\Sigma=d\Omega-\Gamma\wedge\Omega=3dx\wedge dy$. We can then compute the potentials: $\phi=H(e^{-\gamma}\Omega)=H(x^{2}dy-xydx)=0$. Since $d\Gamma=0$, we have $\Sigma^{\prime}=\Sigma$, and $\eta=H(e^{-\gamma}\Sigma)=H(3xdx\wedge dy)=x^{2}dy-xydx=x\Omega$.
For an another choice, let $e^{-\gamma}=x^{2}$. Analogous calculations yield $\gamma=-\ln(x^{2})$, $\Gamma=\frac{-2}{x}dx$, $\Gamma\wedge\Omega=-2dx\wedge dy$, and $\Sigma=\Sigma^{\prime}=4dx\wedge dy$. The potentials in this case are $\phi=0$ and $\eta=x^{3}dy-x^{2}ydx$.
Comparing these two selections of the $e^{-\gamma}$ term demonstrates that the Frobenius decomposition is not fixed; the corresponding terms $\Gamma$ and $\Sigma$ can be adjusted to find a case that is easier to calculate.
@@
$\Omega=Hd\omega=\omega-dH\omega=H(2ydy\wedge dx)=\frac{2}{3}(y^{2}dx-xydy).$ (32)
Note that both terms of the original form $\omega$ contribute to the exact part. This shows that even if a component of a form is not exact, it can still contribute to the exact part of the total form. The potential for the exact term is
@@
We therefore have $e^{\gamma}=\frac{2}{3}xy^{2}$ and $\phi=\ln\left(\frac{x}{y}\right)$, and it is visible that it is a gradient recursive case.
@@ 4 Discussion
The exact part of the work form is associated with the conservative force, a classical result in mechanics. The antiexact part, as noted in [19], can be thought of as the influence of an external agent (or a non-autonomous influence) acting on the system.
Its further decomposition using the Frobenius theorem extracts the integrable part, $e^{\gamma}d\phi$, which can be interpreted in terms of inaccessibility, as in the Carathèodory interpretation of thermodynamics; see [12, 26, 27]. The term $e^{\gamma}d\phi$ represents the force that is aligned along the (integrable) submanifolds, while $e^{\gamma}\eta$ does not define such a geometric structure.
In the gradient recursive case, this is the only term comprising the antiexact part $\Omega$. It can be interpreted as a standard potential-driven force derived from $\phi$, but where the system’s coupling to this potential is modulated by a position-dependent scaling factor $e^{\gamma}$. Physically, it could model phenomena such as a charged particle moving in a magnetic field whose field strength varies in a specific way, or certain types of controlled robotic manipulations.
For the full and recursive case (but not for the gradient recursive case), the term $e^{\gamma}\eta$ is also present. It represents a more significant obstruction to integrability, as it cannot be written as the exterior derivative of a function multiplied by another function. This term represents the ”path-dependent core” of the force, which cannot be eliminated by a simple scaling factor. It can be present in driven systems where the external agent’s input is not derivable from any potential, scaled, or otherwise. Note that this term is present in the full case (when torsion $\Sigma\neq 0$) or in the recursive case (when torsion vanishes, however, when $\Gamma$ has an antiexact part). It therefore measures the torsion and the antiexactness of $\Gamma$.
@@
The operator $d^{\nabla}$ bears a resemblance to the exterior covariant derivative on the trivial line bundle (or fiber set [19, 10]) $U\times\mathbb{R}\rightarrow U$. In the general problem of parallel transport on a bundle, one provides the connection form $\Gamma$ and an external force $\Sigma$ to compute $\Omega$, as was presented in [19, 10]. However, for (6), we obtained the inverse problem: determine $\Gamma$ and $\Sigma$ for a fixed $\Omega$. This is a non-unique problem. Equation (35) also suggests that the antiexact force is a solution of a non-uniform parallel transport problem on a bundle with connection form $\Gamma$ and non-uniformity term $\Sigma$. As we infer in [19], the antiexact part $\Omega$ represents an external influence on the system, and equation (35) can be interpreted as an evolution of this external agent, seen through the force $\Omega$ with which it influences the system. Moreover, the curvature for this external system is given by [^5] In general, the curvature of a connection form $\Gamma$ is given by $\Theta=d\Gamma+\Gamma\wedge\Gamma$. However, for a $\mathbb{R}$-valued form the term $\Gamma\wedge\Gamma$ vanishes. $\Theta=d\Gamma$.
The approach from the paper can also be applied to non-conservative forces, e.g., in hyperelasticity [27], where the work form is an essential construction. The approach presented in the paper can lead to decomposing the work form into smaller components that can be further analyzed. The $\Omega$ represents the non-potential part, and $\eta$ represents the fundamental obstruction to non-potentiality and enables us to quantify it.
@@ 5 Conclusions
A local decomposition for classical forces in a space of arbitrary dimension is proposed. This method employs geometric decomposition on a star-shaped subset of a manifold, distinguishing the exact part, associated with the gradient of a potential, from the antiexact part, which generalizes the curl component of the force. The antiexact part is further resolved using the (non-unique) Frobenius decomposition into an integrable component and a remainder that cannot be represented as the exterior derivative of a function.
@@ Acknowledgments
The authors declare no conflict of interest.
@@
The work was partially funded by the EU Horizon Europe MSCA grant No 101183077. The author would like to thank Michael V. Berry and Pragya Shukla for the enlightening discussion on the nature of curl forces.